# HENASY: Learning to Assemble Scene-Entities for Interpretable Egocentric Video-Language Model

### v2

@@ 2 Related Works
Video-Language Pre-Trained Models. Pre-training VLMs on a large-scale dataset of video-text pairs and deploying them in down-stream tasks has now become a standard practice. Transformer-powered pre-trained VLMs [16, 17, 18, 3, 6, 5, 4, 9] have accomplished superior results on a wide range of tasks, i.e., text-to-video retrieval, action recognition, or events localization. VLMs can be divided into two common categories, i.e., unified- and dual-encoder. The former models [18, 16] fuse multimodal via cross-attention and can be trained with proxy tasks of masked language modeling [19, 20] or masked frame modeling [18, 16]. The latter models [21, 3, 5, 6, 9] employ separate encoders for video and text, trained jointly via contrastive learning [8, 3].
@@ 3 Preliminaries
$\mathcal{L}^{v2t}_{ego}=\frac{1}{\widetilde{B}}\sum_{i\in\widetilde{B}}\log\frac{\sum_{p\in P_{i}}\exp(\hat{\textbf{v}}^{T}\textbf{t}_{p}/\tau)}{\sum_{n\in B}\exp(\hat{\textbf{v}}^{T}\textbf{t}_{n}/\tau)\exp(\hat{\textbf{v}}^{T}\textbf{t}_{n^{\prime}}/\tau)}\quad\text{where $\tau$ denotes a temperature}$ (1)
The objective of text-to-video $\mathcal{L}^{t2v}_{ego}$ is derived from Eq. 1 by inverting $v$ and $t$, and EgoNCE loss is a summation of both directions $\mathcal{L}_{ego}=\mathcal{L}^{v2t}_{ego}+\mathcal{L}^{v2t}_{ego}$.
@@ 4.1 Global Encoder
Global encoder provides global visual information of the entire input video. We adopt the pre-trained TimeSFormer [26] to capture the global visual context of the entire input video. Particularly, we follow the protocol of [3, 5, 4] to decompose the given input video sequence $\mathcal{V}_{i}\in\mathbb{R}^{T\times 3\times H\times W}$ with $T$ RGB frames or resolution $H\times W$ into $T\times K$ non-overlapping patches of resolution $P\times P$, where $K=HW/P^{2}$. Then, every patch is linearly projected by a 2D convolutional layer, forming video patch tokens $\mathbf{z}\in\mathbb{R}^{TK\times D}$ (with $TK=T\times K$) representing embedding features at every temporal and spatial location, where $D$ is hidden dimension. TimeSFormer processes the video patch tokens $\mathbf{z}$ with an additional learnable video tokens $\mathbf{c}\in\mathbb{R}^{1\times D}$ through a stack of divided space-time attention (DST) blocks, which is described as follows: $[\mathbf{c}^{l+1};\mathbf{z}^{l+1}]=\text{DST}([\mathbf{c}^{l};\mathbf{z}^{l}])\quad\text{where $[\cdot;\cdot]$ denotes concatenation operator}$ Please see Sec. A in the appendix for more details of a DST block.
@@ 4.2 Local Entity Encoder
$\mathbf{g}^{l+1}_{boot}=\text{CrossAtt}^{l}(\mathbf{g}^{l}_{boot},\mathbf{z}^{l}),\quad\text{for }l=0,..,S_{1}-1$ (2)
@@
$\mathbf{s}^{l}=\text{TAG}(\mathbf{g}^{l}_{boot},\mathbf{z}^{l})\quad\text{, where $l=S_{1}$}$ (3)
Herein, $\text{TAG}(\mathbf{q},\mathbf{k})$ merges key tokens $\mathbf{k}$ together based on their similarities with query $\mathbf{q}$, while preserving the temporal dimension of key tokens (discussed in detail later). As a result, we obtain new segment tokens $\mathbf{s}^{l}\in\mathbb{R}^{TG\times D}$, which is then utilized as inputs to the entity grouping stage.
Entity Grouping Stage. From this point, local entity encoder is decoupled from the global encoder and is trained to merge these input segments s into complete scene entities e. At this stage, a new set of learnable group tokens $\mathbf{g}^{l}_{entity}\in\mathbb{R}^{E\times D}$ is introduced, which aims to relate segment tokens with similar semantics into an individual scene entity. It is important to note that maintaining consistent temporal dynamics is required at each stage. Therefore, we adopt $S_{2}$ DST blocks [26] to propagate information mutually between learnable group tokens and segment tokens: $[\mathbf{g}^{l+1}_{entity};\mathbf{s}^{l+1}]=\text{DST}([\mathbf{g}^{l}_{entity};\mathbf{s}^{l}])$, for $l=S_{1},..,S_{1}+S_{2}-1$.
After the final layer, segment tokens $\mathbf{s}^{l}$ are grouped to generate intermediate entity tokens, i.e., $\hat{\mathbf{e}}^{l}=\text{TAG}(\textbf{g}^{l}_{entity},\textbf{s}^{l})\in\mathbb{R}^{TE\times D},\text{ where }l=S_{1}+S_{2}$. Then, to enable interactions between scene entities and across temporal dimension, we apply a stack of $S_{3}$ DST blocks to all entity tokens: $\hat{\mathbf{e}}^{l+1}=\text{DST}(\hat{\mathbf{e}}^{l})$, for $l=S_{1}+S_{2},..,S_{1}+S_{2}+S_{3}-1$.
We observed that segment tokens $\mathbf{s}$ and entity tokens $\mathbf{e}$ facilitate temporal consistencies within the temporal attention of a DST block. Unlike in TSF [26], where tokens are spatially limited within a patch, these tokens can evolve freely across frames, enhancing the flexibility of the time-attention mechanism in a DST block. To obtain spatio-temporal entity embeddings, we apply temporally average pooling on entity tokens of the final layer: $\mathbf{e}=\text{AvgPool}(\hat{\mathbf{e}}^{l})$, where $\mathbf{e}\in\mathbb{R}^{E\times D}$ and $l=S_{1}+S_{2}+S_{3}$.
@@
$\mathbf{o}=\text{TAG}(\mathbf{g}_{q},\mathbf{i})\text{ ; }\quad\mathbf{o}_{t,i}=\text{TAG}_{t,i}(\mathbf{g}_{q},\mathbf{i})=({\mathbf{g}_{q}})_{i}+\frac{\sum^{I}_{j=1}\tilde{\mathbf{A}}_{t,i,j}\mathbf{i}_{t,j}}{\sum^{I}_{j=1}\tilde{\mathbf{A}}_{t,i,j}}$ (4)
Afterwards, we re-shape $\mathbf{o}$ to $\mathbb{R}^{TQ\times D}$ as output of TAG. The computation of similarity A and assignment $\tilde{\mathbf{A}}$ matrices, along with deriving saliency maps of assignments between video patches and entities for interpretability, is detailed in Sec. B in appendix.
@@ 4.3 Entity-Aware Decoder
$\begin{split}\tilde{\mathbf{e}}^{b_{dec}}&=\text{CrossAtt}(\mathbf{e}^{b_{dec}},\mathbf{z}^{l})\\ \mathbf{e}^{b_{dec}+1}&=\text{MLP}\big{(}\text{SelfAtt}(\tilde{\mathbf{e}}^{b_{dec}})\big{)}\end{split}$ (5)
@@
$\mathbf{v}=\text{AvgPool}(\mathbf{e}^{b_{dec}})$ (6)
@@ 4.4 Multi-grained Contrastive Learning
Projection Loss operates on each individual frame of the input video, utilizing an external object detector [27] to identify bounding boxes $b=\{b_{i}\in\mathbb{R}^{4}\}_{i=1}^{N_{b}}$ of scene entities. Let $\mathbf{m}=\{\mathbf{m}_{i}\in(0,1)^{H\times W}\}_{i=1}^{E}$ be the predicted foreground probability maps of scene entities, Hungarian matching pairs each detected box $b_{i}$ with the predicted mask $m_{i}$ having the highest IoU.
@@ 5.1 Training and Implementation Details
Training. HENASY is trained on EgoClip [3], which contains 3.8M clip-narration pairs covering a sub-set of 2,927 video hours from Ego4D [1]. For each video clip, we uniformly sample 4 frames. We employ the pre-extracted narration’s nouns and pre-detected hand and object bounding boxes from [4] for NEC loss and projection loss, respectively. For verb phrases, we employ LLama-2 [28] with a prompt as discussed in Sec.C. The loss weights in Eq. 10 are set as: $\lambda_{1}=0.5,\lambda_{2}=0.5,\lambda_{3}=1.0$. We train HENASY on two A6000 GPUs, in 5 epochs with AdamW optimizer [31] at fixed learning rate of $53e-$ 5, and with batch size of 128. We initialize global encoder and text encoder with pretrained model provided from [5], but freeze them in the entire training process.
@@ 5.2 Benchmarks and Evaluation Protocols
Table 1: Comparison on the zero-shot transfer over EgoMCQ, EK100-MIR, EK100-CLS, and EGTEA.
@@ V-T T-V Avg V-T T-V Avg Acc Acc Acc Acc
5.3 Main Results
Table 2: Comparison on the visual & Textual representation over Ego-NLQ and EgoMQ.
@@ 5.4 Ablation Studies
Impact of Entity-Aware Decoder: We assess the effect of entity-aware (EA) decoder on zero-shot tasks in the first two rows of Table 4. In the first experiment, we omit the proposed decoder and simply operate entity tokens through an average pooling to obtain video representations, the performance drops significantly (2% across benchmarks) compared to using the proposed decoder as shown in the second experiment. These ablation experiments show that modeling interactions between global features and entity embeddings plays an important role in our model, and the proposed design of entity-aware decoder is beneficial to the overall performance.
Impact of Bootstrapping Stage: Next, we report the effect of bootstrapping stage in the third row of Table 4, where we remove bootstrapping stage by directly processing video patch tokens. The performance degrades by 1% across all benchmarks, showing the effectiveness of this design choice.
Table 3: Ablation results on model design
Model designs EgoMCQ EK100-MIR EK100-CLS
Inter Intra Avg Avg Top-1 Top-5
mAP nDCG Acc Acc
w/o EA Decoder 87.6 47.9 18.8 25.5 6.7 18.1
w/ EA Decoder 93.3 59.1 30.4 32.8 18.0 36.3
w/o Bootstrapping 92.6 59.2 31.1 32.6 19.2 37.9
complete settings 94.1 61.3 31.3 33.2 19.5 38.2
@@
Table 5: Ablation on computational complexity and memory cost
@@
GFLOPs (per clip) 530M 599M
Number of Parameters 216M 291M
GPU Memory (train) 38GB 42GB
GPU Memory (inference) 4.4GB 4.8GB
@@ 6 Conclusions
Appendix

### v4

@@ Abstract
Project page: https://uark-aicv.github.io/HENASY
@@ 2 Related Works
Video-Language Pre-Trained Models. Pre-training VLMs on a large-scale dataset of video-text pairs and deploying them in down-stream tasks has now become a standard practice. Transformer-powered pre-trained VLMs [16, 17, 18, 19, 3, 6, 5, 4, 9] have accomplished superior results on a wide range of tasks, i.e., text-to-video retrieval, action recognition, or events localization. VLMs can be divided into two common categories, i.e., unified- and dual-encoder. The former models [16, 18, 20] fuse multimodal via cross-attention and can be trained with proxy tasks of masked language modeling [21, 22] or masked frame modeling [18, 16]. The latter models [23, 3, 5, 6, 9] employ separate encoders for video and text, trained jointly via contrastive learning [8, 3].
@@
Interpretable Video Representation. There have been efforts to enhance the interpretability of video representations [24, 25, 26, 27], these typically involved factorizing videos into entities and environmental contexts, and utilizing adaptive attention mechanisms [24] to selectively focus on relevant entities. Such mechanisms enable interpretation through their selection module, highlighting only the primary entities contributing to the final predictions. However, these approaches often require entities to be pre-detected by external models, which is restricted to a predefined set of objects. Additionally, these works have focused on models specifically designed for individual tasks, e.g., temporal action detection, video dense captioning. In contrast, our work introduces an end-to-end method capable of learning to form entities without an off-the-shelf detector. Moreover, we aim to develop a robust video-language model (VLM) that is versatile across a variety of tasks.
@@ 3 Preliminaries
$\mathcal{L}^{v2t}_{ego}=\frac{1}{\widetilde{B}}\sum_{i\in\widetilde{B}}\log\frac{\sum_{p\in P_{i}}\exp(\hat{\textbf{v}}^{T}\hat{\textbf{t}}_{p}/\tau)}{\sum_{n\in B}\exp(\hat{\textbf{v}}^{T}\hat{\textbf{t}}_{n}/\tau)+\exp(\hat{\textbf{v}}^{T}\hat{\textbf{t}}_{n^{\prime}}/\tau)}\quad\text{where $\tau$ denotes a temperature}$ (1)
The objective of text-to-video $\mathcal{L}^{t2v}_{ego}$ is derived from Eq. 1 by inverting $v$ and $t$, and EgoNCE loss is a summation of both directions $\mathcal{L}_{ego}=\mathcal{L}^{v2t}_{ego}+\mathcal{L}^{t2v}_{ego}$.
@@ 4.1 Global Encoder
Global encoder provides global visual information of the entire input video. We adopt the pre-trained TimeSFormer [32] to capture the global visual context of the entire input video. Particularly, we follow the protocol of [3, 5, 4] to decompose the given input video sequence $\mathcal{V}_{i}\in\mathbb{R}^{T\times 3\times H\times W}$ with $T$ RGB frames or resolution $H\times W$ into $T\times K$ non-overlapping patches of resolution $P\times P$, where $K=HW/P^{2}$. Then, every patch is linearly projected by a 2D convolutional layer, forming video patch tokens $\mathbf{z}\in\mathbb{R}^{TK\times D}$ (with $TK=T\times K$) representing embedding features at every temporal and spatial location, where $D$ is hidden dimension. TimeSFormer processes the video patch tokens $\mathbf{z}$ with an additional learnable video tokens $\mathbf{c}\in\mathbb{R}^{1\times D}$ through a stack of divided space-time attention (DST) blocks, which is described as follows: $[\mathbf{c}^{l+1};\mathbf{z}^{l+1}]=\text{DST}([\mathbf{c}^{l};\mathbf{z}^{l}])\quad\text{where $[\cdot;\cdot]$ denotes concatenation operator}$. Please see Appendix A for more details on the computational process of a DST block.
@@ 4.2 Local Entity Encoder
$\mathbf{g}^{l+1}_{boot}=\texttt{CrossAtt}^{l}(\mathbf{g}^{l}_{boot},\mathbf{z}^{l}),\quad\text{for }l=0,..,S_{1}-1$ (2)
@@
$\mathbf{s}^{l}=\texttt{TAG}(\mathbf{g}^{l}_{boot},\mathbf{z}^{l})\quad\text{, where $l=S_{1}$}$ (3)
Herein, $\texttt{TAG}(\mathbf{q},\mathbf{k})$ merges key tokens $\mathbf{k}$ together based on their similarities with query $\mathbf{q}$, while preserving the temporal dimension of key tokens (discussed in detail later). As a result, we obtain new segment tokens $\mathbf{s}^{l}\in\mathbb{R}^{TG\times D}$, which is then utilized as inputs to the entity grouping stage.
Entity Grouping Stage. From this point, local entity encoder is decoupled from the global encoder and is trained to merge these input segments s into complete scene entities e. At this stage, a new set of learnable group tokens $\mathbf{g}^{l}_{entity}\in\mathbb{R}^{E\times D}$ is introduced, which aims to relate segment tokens with similar semantics into an individual scene entity. It is important to note that maintaining consistent temporal dynamics is required at each stage. Therefore, we adopt $S_{2}$ DST blocks [32] to propagate information mutually between learnable group tokens and segment tokens: $[\mathbf{g}^{l+1}_{entity};\mathbf{s}^{l+1}]=\texttt{DST}([\mathbf{g}^{l}_{entity};\mathbf{s}^{l}])$, for $l=S_{1},..,S_{1}+S_{2}-1$.
After the final layer, segment tokens $\mathbf{s}^{l}$ are grouped to generate intermediate entity tokens, i.e., $\hat{\mathbf{e}}^{l}=\texttt{TAG}(\textbf{g}^{l}_{entity},\textbf{s}^{l})\in\mathbb{R}^{TE\times D},\text{ where }l=S_{1}+S_{2}$. Then, to enable interactions between scene entities and across temporal dimension, we apply a stack of $S_{3}$ DST blocks to all entity tokens: $\hat{\mathbf{e}}^{l+1}=\texttt{DST}(\hat{\mathbf{e}}^{l})$, for $l=S_{1}+S_{2},..,S_{1}+S_{2}+S_{3}-1$.
We observed that segment tokens $\mathbf{s}$ and entity tokens $\mathbf{e}$ facilitate temporal consistencies within the temporal attention of a DST block. Unlike in TSF [32], where tokens are spatially limited within a patch, these tokens can evolve freely across frames, enhancing the flexibility of the time-attention mechanism in a DST block. To obtain spatio-temporal entity embeddings, we apply temporally average pooling on entity tokens of the final layer: $\mathbf{e}=\texttt{AvgPool}(\hat{\mathbf{e}}^{l})$, where $\mathbf{e}\in\mathbb{R}^{E\times D}$ and $l=S_{1}+S_{2}+S_{3}$.
@@
$\mathbf{o}=\texttt{TAG}(\mathbf{g}_{q},\mathbf{i})\text{ ; }\quad\mathbf{o}_{t,i}=\texttt{TAG}_{t,i}(\mathbf{g}_{q},\mathbf{i})=({\mathbf{g}_{q}})_{i}+\frac{\sum^{I}_{j=1}\tilde{\mathbf{A}}_{t,i,j}\mathbf{i}_{t,j}}{\sum^{I}_{j=1}\tilde{\mathbf{A}}_{t,i,j}}$ (4)
Afterwards, we re-shape $\mathbf{o}$ to $\mathbb{R}^{TQ\times D}$ as output of TAG.
The computation of similarity matrix A and assignment matrix $\tilde{\mathbf{A}}$ are detailed in Appendix B.1. Additionally, the derivation of saliency maps for interpreting assignments between video patches and entities is explained Appendix B.2.
@@ 4.3 Entity-Aware Decoder
$\begin{split}\tilde{\mathbf{e}}^{b_{dec}}&=\texttt{CrossAtt}(\mathbf{e}^{b_{dec}},\mathbf{z}^{l})\\ \mathbf{e}^{b_{dec}+1}&=\texttt{MLP}\big{(}\texttt{SelfAtt}(\tilde{\mathbf{e}}^{b_{dec}})\big{)}\end{split}$ (5)
@@
$\mathbf{v}=\texttt{AvgPool}(\mathbf{e}^{b_{dec}})$ (6)
@@ 4.4 Multi-grained Contrastive Learning
Projection Loss operates on each individual frame of the input video, utilizing an external object detector [33] to identify bounding boxes $b=\{b_{i}\in\mathbb{R}^{4}\}_{i=1}^{N_{b}}$ of scene entities. Let $\mathbf{m}=\{\mathbf{m}_{i}\in(0,1)^{H\times W}\}_{i=1}^{E}$ be the predicted saliency maps of scene entities (see Appendix B.2 for saliency maps formulation), Hungarian matching pairs each detected box $b_{i}$ with the predicted mask $\mathbf{m}_{i}$ having the highest IoU.
@@ 5.1 Training and Implementation Details
Training. HENASY is trained on EgoClip [3], which contains 3.8M clip-narration pairs covering a sub-set of 2,927 video hours from Ego4D [1]. For each video clip, we uniformly sample 4 frames. We employ the pre-extracted narration’s nouns and pre-detected hand and object bounding boxes from [4] for NEC loss and projection loss, respectively. For verb phrases, we employ Llama-2 [34] with a prompt as discussed in Appendix C. The loss weights in Eq. 10 are set as: $\lambda_{1}=0.5,\lambda_{2}=0.5,\lambda_{3}=1.0$. We train HENASY on two A6000 GPUs, in 5 epochs with AdamW optimizer [37] at fixed learning rate of $53e-$ 5, and with batch size of 128. We initialize global encoder and text encoder with pretrained model provided from [5], but freeze them in the entire training process.
@@ 5.2 Benchmarks and Evaluation Protocols
Table 1: Comparison on the zero-shot transfer over EgoMCQ, EK100-MIR, EK100-CLS, and EGTEA. HelpingHands* refers to our re-produced results with TSF-B backbone using provided source code [4] .
@@ V-T T-V Avg V-T T-V Avg Acc Acc Acc Acc
Table 2: Comparison on the visual & textual representation over EgoNLQ and EgoMQ. Grey indicates result we obtained using provided pre-trained checkpoint that.
@@ 5.4 Ablation Studies
Impact of Entity-Aware Decoder: We evaluate the influence of the entity-aware (EA) decoder on zero-shot tasks in the first two rows of Table 5. In the first experiment (labeled as ‘w/ avg. pool’), the proposed EA decoder is omitted, and entity tokens are merely processed using average pooling to generate the video representation. This configuration results in a notable decline in performance across all benchmarks. In the second experiment (labeled as ‘w/ SA Dec.’), video patch tokens are combined with entity tokens and then input into a self-attention decoder, which has the same dimensions as our EA decoder. This setup leads to a performance decrease of approximately 2% across benchmarks compared to our proposed decoder (‘complete settings’).
These ablation studies show that modeling interactions between global features and entity embeddings plays a critical role, and the proposed design of entity-aware decoder significantly enhances overall model performance.
Table 4: Ablation results on model design.
Model designs EgoMCQ EK100-MIR EK100-CLS
Inter Intra Avg Avg Top-1 Top-5
mAP nDCG Acc Acc
w/ avg. pool 87.6 47.9 18.8 25.5 6.7 18.1
w/ SA dec. 93.3 59.1 30.4 32.8 18.0 36.3
w/o bootstrap 92.6 59.2 31.1 32.6 19.2 37.9
complete settings 94.1 61.3 31.3 33.2 19.5 38.2
Table 5: Comparison on computational complexity and memory cost.
@@
GFLOPs per video clip (Million) 530 599
Number of Parameters (Million) 216 291
Train GPU Memory (GB) 38 42
Inference GPU Memory (GB) 4.4 4.8
@@
Impact of Bootstrapping Stage: We report the effect of bootstrapping stage in the third row of Table 5 (labeled as w/o bootstrap), where we remove bootstrapping stage by directly processing video patch tokens. The performance degrades by 1% across all benchmarks, showing the effectiveness of this design choice.
@@ 6 Conclusions
Acknowledgements
The authors gratefully acknowledge funding supports from U.S. National Science Foundation (NSF) under Award No. OIA-1946391 and NSF EFRI BRAID 2223793.
@@ Appendix C Verb Phrase Generation
Appendix D Quantitative evaluation on visual grounding
We conducted a rigorous quantitative analysis on Ego4D dataset to compare with the SOTA model of HelpingHands, as they only provide visual grounding results on this datasets. We create semantic segmentation labels by ourselves for 200 videos. The results are reported under mIoU metrics between visual grounding prediction with corresponding groundtruth, showing our model’s superior visual grounding capability compared to HelpingHands:
Model mIoU
HelpingHands [4] 22.73%
Ours 41.06%
Table 6: Comparison with HelpingHands on visual grounding task.
Discussion: Despite being the SOTA in egocentric tasks and visual grounding task, HelpingHands [4] only provides coarse bounding boxes of objects, leading to much lower mIoU scores due to inadequate coverage of the target masks. In contrast, our model employs segmentation masks that closely align with the ground truth, resulting in higher mIoU scores, demonstrating superior visual grounding capabilities.